OpenStudy (anonymous):

What is the value of (3+5^2)divided by 4 -(x+1) when x=7?

OpenStudy (mertsj):

\[\frac{3+5^{2}}{4-(x+1)}=\frac{28}{3-x}=\frac{28}{3-7}=\frac{28}{-4}=-7\]